OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1?

OpenStudy (anonymous):

if y=mx +c be parallel to y=kx+d then m=k in other words parallel lines have equal slope

OpenStudy (anonymous):

first find the slope of 4x + y = -1 slope of ax +by=c is given by -(coefficient of x)/(coefficient of y) and this will also be the slope of the required parallel line with the given one

OpenStudy (anonymous):

im confused

OpenStudy (e.mccormick):

Know the standard form \(y=mx+b\) where m is the slope?  Well, parallel lines have the same slope. So put it in \(y=mx+b\) form, then m is your answer.

OpenStudy (anonymous):

m=2?

OpenStudy (anonymous):

or m=-4?

OpenStudy (e.mccormick):

Yes! -4!

OpenStudy (e.mccormick):

Not sure where that 2 came from. hehe.

OpenStudy (anonymous):

Lol me either im a little slow

OpenStudy (e.mccormick):

Well, you kept with it and got the right answer. That is what matters!